# Toward Efficient and Robust Behavior Models for Multi-Agent Driving Simulation

### v3

@@ I INTRODUCTION
The second requirement is achieving high execution speed to enable efficient scaling. While prior works, e. g., [14, 16, 27, 10], have focused on accelerating the simulation backbone, our focus is on optimizing the executed behavior model. Commonly, agent-centric observations (see Figure 1a) are used, e. g., [20, 16, 10, 9], where nearby instances, such as agents and static map elements, are represented from each agent’s perspective. However, generating and encoding these local observations for every agent at each simulation step imposes a significant computational burden. Alternatively, scene-centric observations (see Figure 1b) represent all instances in a global coordinate frame, e. g., [32, 23]. While this offers efficient encoding, it comes at the potential sacrifice of pose-invariance, requiring the model to implicitly learn variations in agent positions and orientations.
Instead, to improve the efficiency of the behavior model, we employ instance-centric observations (see Figure 1c), where each instance, i. e., an individual agents or map elements, is represented in its own local coordinate frame. Features are encoded using shared context encoders, allowing the encoded instance tokens to be processed across multiple agents. This approach models relationships between tokens in a symmetric manner, independent of any global coordinate frame. As a result, static instances need to be encoded only once, and dynamic instances only once per simulation step, rather than from every agent’s perspective at each step.
Our Contributions are four-fold: (1) Our main contribution is the introduction of an instance-centric scene representation for behavior modeling, which enables efficient, viewpoint-invariant encoding of the scene. Instance tokens are shared across all agents. Static instances need to be encoded only once, enabling the reuse of their tokens in subsequent simulation steps. (2) We employ a query-centric symmetric context encoder with relative positional encodings between pairwise instance-centric frames, and evaluate its performance across different model design choices. (3) We propose an adaptive reward transformation that automatically balances robustness and realism during AIRL training. (4) We demonstrate that our approach yields realistic behavior models, outperforming diverse baselines in robustness and scalability across two automated driving datasets.
@@ II Related Work
Realistic Behavior Simulation: Early approaches to simulating vehicle behavior include log-replay [4, 14], where trajectories recorded in the real world are simply replayed, rule-based models [17, 18, 14], as well as Behavior Cloning (BC) [22, 3], an open-loop IL method, where a mapping from observations to actions is learned from a dataset using supervised learning. Due to its simplicity and the availability of large-scale datasets, the latter remains widely used. Each approach, however, has its limitations: Replayed trajectories are non-reactive, thus prohibiting interactions with a simulated ego vehicle. Rule-based models often fail to capture the complexity of real-world behaviors. BC performs poorly in rare or unseen situations (e. g., near-collisions or very aggressive merges). Even worse, when deploying such a model in a closed-loop simulation, small deviations from expert driving accumulate, leading the model into unfamiliar states - known as covariate shift [29].
For learning more robust behavior models, RL is commonly employed [6, 9, 10, 19, 2]. Since realistically simulating traffic situations is inherently a multi-agent problem, self-play RL has emerged as an effective solution. In [19], all agents are controlled by a single shared policy, which is trained jointly through self-play. To enable applying the model to a diverse set of traffic situations, a flexible, graph-based input representation with vectorized map features is adopted. Similarly, in [10] and [9], self-play RL has been applied on a large scale, demonstrating that robust and naturalistic behavior emerges from it. The resulting behavior model generalizes well, even in out-of-distribution scenarios. Notably, in [10], the behavior model is trained in randomly initialized traffic situations, driving more than $1.6$ billion km during training. The resulting behavior model is zero-shot evaluated on three autonomous driving benchmarks, achieving state-of-the-art performance without ever seeing human data during training.
@@
All of the discussed approaches to behavior modeling employ agent-centric observations, where each observation is expressed relative to the simulated agent’s frame of reference. While straightforward, this becomes computationally expensive when simulating multiple agents, as the complexity scales linearly with the number of agents and quadratically with the number of pairwise interactions between agents.
@@ III Method
As illustrated in Figure 2, simulations evolve recursively. First, an observation model generates a description of the traffic situation, either in an agent-centric or instance-centric manner. Next, a learned behavior model maps the observation to the actions that the agent should execute, i. e., its acceleration and steering angle. Finally, the selected actions are executed via a kinematics model, updating the agent’s state. Repeating these steps for all agents at fine-grained time intervals enables continuous traffic simulation. Our in-house simulation backbone is implemented in C++, thus achieving a low runtime. Figure 3 shows an example situation.
This sequential decision-making process is formulated as a Partially-Observable Markov Decision Process (POMDP), characterized by the tuple $(S,O,A,T,R,\Omega,\gamma)$. In this framework, the agent does not have direct access to the true state $s\in S$ of the environment. Instead, it receives a noisy observation $o\in O$, provided by the observation model $\Omega:S\to O$. Upon executing an action $a\in A$, the state of the environment is updated stochastically according to the transition probability density $T:S\times A\times S\rightarrow\left[0,\infty\right[$. In addition, the agent receives a numerical reward defined by the reward function $R:S\times A\to\mathbb{R}$, as well as a new observation of the updated environment state. The discount factor $\gamma\in[0,1[$ balances the trade-off between immediate and future rewards. The solution to a POMDP is the optimal policy $\pi^{*}:O\times A\to[0,\infty[$, which maps observations to action distributions, maximizing the expected cumulative reward $J(\pi)=\mathop{\mathbb{E}}_{\pi}\left[\sum_{k=0}^{\infty}\gamma^{k}r_{k}\right]$, where $r_{k}=R(s_{k},a_{k})$.
@@ III-B Adversarial Inverse Reinforcement Learning
RL requires defining a reward function that accurately captures the incentive structure of real-world driving. Instead of manually defining such a reward function, we use AIRL [12] to reconstruct a surrogate reward signal from real data, given as $\mathcal{D}=\{(o_{1},a_{1}),(o_{2},a_{2}),\dots\}$. In AIRL, an additional discriminator model $D_{\phi}$ is trained to distinguish generated from real samples, outputting the probability $D_{\phi}(o,a)\in\left[0,1\right]$ for the observation-action pair being real, i. e., stemming from $\mathcal{D}$. The policy is trained via RL using the surrogate reward
@@
which guides the policy to taking realistic actions. Following [20], to smooth the decision boundary during discriminator training, we add random noise to the actions in $\mathcal{D}$, matching the standard deviation of the policy.
Figure 2: Single simulation step: The behavior model maps observations to actions, which are then executed via a kinematic bicycle model.
Figure 3: Illustration of an example situation using instance-centric observations. Simulated and corresponding ground-truth vehicles are depicted as solid and outlined rectangles, respectively.
Figure 4: Illustration of the proposed instance-centric behavior model mapping observations to actions. Instance encoders convert observations into latent tokens. These tokens are then augmented with positional encodings relative to the target agent before being passed through multiple layers of our refinement module. Lastly, each refined actor token is decoded into its corresponding actions. Static map tokens can be reused across simulation steps.
@@
However, using a fixed $c$ across experiments, e. g., when comparing different discriminator architectures or scene representations, may result in varying expected values of $\tilde{r}$, thereby influencing the agent’s incentive to survive. For instance, a weaker discriminator tends to learn a less distinct decision boundary, yielding higher surrogate rewards, whereas a stronger discriminator learns sharper boundaries, producing lower surrogate rewards. Consequently, policies trained with stronger discriminators have less incentive to avoid collisions and stay on the road, despite in both cases the optimal behavior w. r. t. the discriminator is to match the distribution of the real data $\mathcal{D}$.
To decouple the agent’s survival incentive from the discriminator’s performance, we propose an adaptive reward offset
@@
where $\tilde{r}_{\mathrm{mean}}(n)$ is the average surrogate reward for the generated samples in epoch $n$, and $\tilde{r}_{\mathrm{target}}$ is a hyperparameter, specifying the desired average reward. In other words, our dynamic reward offset can be interpreted as an additional reward term that encourages the policy to survive longer, especially when the discriminator is confident that its samples are generated. This adaptive approach keeps the expected value of $\tilde{r}$ consistent, thus enabling fair comparisons across experiments, while still allowing the discriminator to guide the policy toward more realistic behavior by computing the appropriate offset to $\tilde{r}_{\mathrm{target}}$.
@@ III-C Instance-Centric Observations
For static map elements, such as pedestrian crossings or lane boundaries, we use a vectorized representation [13]. Each map element is approximated by a polyline with a maximum total length of $10\text{\,}\mathrm{m}$. A polyline is composed of multiple vectors, with each being described by a set of features $\mathbf{v}=\left[\mathbf{v}_{\mathrm{start}},\mathbf{v}_{\mathrm{end}},\mathbf{v}_{\mathrm{type}}\right]^{\intercal}$, including the vector’s start and end point, as well as a one-hot encoding for the polyline type. For polylines, the origin of the instance-centric coordinate frame is given as the mean of all polyline points, with the x-axis being aligned to the tangential direction of the polyline.
The local coordinate frame of an agent is determined by its last position and moving direction. Agents are described by their size (width and length), velocity, current speed limit, and a binary indicator for VRUs: $\mathbf{x}=\left[\mathbf{x}_{\mathrm{size}},x_{\mathrm{vel}},x_{\mathrm{limit}},x_{\mathrm{vru}}\right]^{\intercal}$. Note that, compared to existing works utilizing agent-centric observations, here the agent’s position and orientation are not included in the agent features, but are now captured by the position and orientation of its local coordinate frame.
@@ III-D Behavior Model Details
followed by an element-wise max-pooling operation over $\{\mathbf{v}_{1},\mathbf{v}_{2},\dots,\mathbf{v}_{L}\}$ to produce the final polyline token $\mathbf{z}^{\mathrm{(map)}}$. We implement $g_{\mathrm{enc}}(\cdot)$ as an multi-layer perceptron (MLP), $f_{\mathrm{agg}}(\cdot)$ as element-wise max-pooling, and $f_{\mathrm{rel}}(\cdot)$ as concatenation. Conversely, each agent feature vector $\mathbf{x}$ is encoded into the same latent space using another MLP, producing the agent token $\mathbf{z}^{\mathrm{(agent)}}=\mathrm{MLP}(\mathbf{x})$.
Note that with agent-centric observations, each instance would need to be encoded $N_{a}\times H$ times — once per agent per timestep, where $H$ is the number of simulation steps. In contrast, with our instance-centric representation, the polyline tokens $Z^{\mathrm{(map)}}=\{\mathbf{z}^{\mathrm{(map)}}_{i}\mid i=1,\cdots,N_{p}\}$ remain unchanged throughout the rollout. This enables reusing the tokens in subsequent simulation steps, reducing the number of encoding a polyline-instance from $N_{a}\times H$ times to just one. However, agent states are updated during the simulation rollouts, requiring agent tokens $Z^{\mathrm{(agent)}}=\{\mathbf{z}^{\mathrm{(agent)}}_{i}\mid i=1,\cdots,N_{a}\}$ to be regenerated every simulation step. Nonetheless, since agents are encoded in a viewpoint-invariant manner, they only need to be encoded $H$ times instead of $N_{a}\times H$ times.
@@
For that, we define the relative positional encoding between a target agent $i\in\{0,\cdots,N_{a}\}$ and any of the combined instance tokens $Z=\left[Z^{\mathrm{(agent)}},\;Z^{\mathrm{(map)}}\right]$, indexed by $j$, as $\mathbf{c}_{i\rightarrow j}=\mathbf{r}_{i\rightarrow j}\oplus\mathds{1}^{(\mathrm{agent})}_{j}\oplus\mathds{1}^{(\mathrm{route})}_{j}$, where $\oplus$ denotes concatenation. The binary indicators $\mathds{1}^{(\mathrm{agent})}_{j}$ and $\mathds{1}^{(\mathrm{route})}_{j}$ specify whether instance $j$ is an agent, or a map polyline that is part of the route that agent $i$ should follow, respectively. These indicators are appended here rather than to the vector features $\mathbf{v}$, as each agent may follow a different route, making a global encoding at the vector level infeasible. Finally, following the approach proposed in [24], we compute pairwise encodings by applying a feature-wise linear transformation:
@@
Decoder: Finally, the decoder processes the refined instance token $\mathbf{z}_{i}^{(K)}$ from the final Perceiver layer. In the behavior model, it outputs the mean and standard deviation of the next acceleration and steering angle. In the discriminator, which evaluates observation-action pairs, the action is appended to $\mathbf{z}_{i}^{(K)}$ to then compute the classification score $D_{\phi}$.
We use three layers of the message passing mechanism defined in (5) in our polyline encoder and $16$ channels per head in (7). Unless specified otherwise, each MLP in the network consists of a linear layer, followed by layer normalization, a ReLU activation, and a second linear layer.

### v5

@@ I INTRODUCTION
The second requirement is achieving high execution speed to enable efficient scaling. While prior works, e. g., [14, 16, 27, 10], have focused on accelerating the simulation backbone, our focus is on optimizing the executed behavior model.
Commonly, agent-centric observations (see Figure 1a) are used, e. g., [20, 16, 10, 9], where nearby instances, such as agents and static map elements, are represented from each agent’s perspective. However, generating and encoding these local observations for every agent at each simulation step imposes a significant computational burden. Alternatively, scene-centric observations (see Figure 1b) represent all instances in a global coordinate frame, e. g., [32, 23]. While this offers efficient encoding, it comes at the potential sacrifice of pose-invariance, requiring the model to implicitly learn variations in agent positions and orientations.
Instead, to improve the efficiency of the behavior model, we employ instance-centric observations (see Figure 1c), where each instance, i. e., an individual agent or map element, is represented in its own local coordinate frame. Features are encoded using shared context encoders, allowing the encoded instance tokens to be processed across multiple agents. This approach models relationships between tokens in a symmetric manner, independent of any global coordinate frame. As a result, static instances need to be encoded only once, and dynamic instances only once per simulation step, rather than from every agent’s perspective at each step.
Our Contributions are four-fold: (1) Our main contribution is the introduction of an instance-centric scene representation for behavior modeling, which enables efficient, viewpoint-invariant encoding of the scene. Instance tokens are shared across all agents. Static instances need to be encoded only once, enabling the reuse of their tokens in subsequent simulation steps. (2) We employ a query-centric symmetric context encoder with relative positional encodings between pairwise instance-centric frames, and evaluate its performance across different model design choices. (3) We propose an adaptive reward transformation that automatically balances robustness and realism during AIRL training. (4) We demonstrate that our approach yields realistic behavior models, outperforming a diverse set of baselines in robustness and scalability across two automated driving datasets.
@@ II Related Work
Realistic Behavior Simulation: Early approaches to simulating vehicle behavior include log-replay [4, 14], where trajectories recorded in the real world are simply replayed, rule-based models [17, 18, 14], as well as Behavior Cloning (BC) [22, 3], an open-loop IL method, where a mapping from observations to actions is learned from a dataset using supervised learning. Due to its simplicity and the availability of large-scale datasets, the latter remains widely used. Each approach, however, has its limitations: Replayed trajectories are non-reactive, thus prohibiting interactions with a simulated ego vehicle. Rule-based models often fail to capture the complexity of real-world behaviors. BC performs poorly in rare or unseen scenarios (e. g., near-collisions or very aggressive merges). Even worse, when deploying such a model in a closed-loop simulation, small deviations from expert driving accumulate, leading the model into unfamiliar states - known as covariate shift [29].
For learning more robust behavior models, RL is commonly employed [6, 9, 10, 19, 2]. Since realistically simulating traffic scenarios is inherently a multi-agent problem, self-play RL has emerged as an effective solution. In [19], all agents are controlled by a single shared policy, which is trained jointly through self-play. To enable applying the model to a diverse set of traffic scenarios, a flexible, graph-based input representation with vectorized map features is adopted. Similarly, in [10] and [9], self-play RL has been applied on a large scale, demonstrating that robust and naturalistic behavior emerges from it. The resulting behavior model generalizes well, even in out-of-distribution scenarios. Notably, in [10], the behavior model is trained in randomly initialized traffic scenarios, driving more than $1.6$ billion km during training. The resulting behavior model is zero-shot evaluated on three autonomous driving benchmarks, achieving state-of-the-art performance without ever seeing human data during training.
@@
All the discussed approaches to behavior modeling employ agent-centric observations, where each observation is expressed relative to the simulated agent’s frame of reference. While straightforward, this becomes computationally expensive when simulating multiple agents, as the complexity scales linearly with the number of agents and quadratically with the number of pairwise interactions between agents.
@@ III Method
As illustrated in Figure 2, simulations evolve recursively. First, an observation model encodes the traffic situation, either in an agent-centric or instance-centric manner. Next, a learned behavior model maps the observation to actions, i. e., acceleration and steering angle for vehicles, and acceleration and heading rate for pedestrians. Finally, these actions are executed via a kinematics model, updating the agent’s state. Repeating these steps for all agents at fine-grained time intervals yields continuous traffic simulation. Our in-house simulation backbone is implemented in C++, thus achieving a low runtime. Figure 3 shows an example scenario.
This sequential decision-making process is formulated as a Partially-Observable Markov Decision Process (POMDP), characterized by the tuple $(S,O,A,T,R,\Omega,\gamma)$. In this framework, the agent does not have direct access to the true state $s\in S$ of the environment. Instead, it receives a noisy observation $o\in O$, provided by the observation model $\Omega:S\to O$. Upon executing an action $a\in A$, the state of the environment is updated stochastically according to the transition probability density $T:S\times A\times S\rightarrow\left[0,\infty\right)$. In addition, the agent receives a numerical reward defined by the reward function $R:S\times A\to\mathbb{R}$, as well as a new observation of the updated environment state. The discount factor $\gamma\in\left[0,1\right)$ balances the trade-off between immediate and future rewards. The solution to a POMDP is the optimal policy $\pi^{*}:O\times A\to\left[0,\infty\right)$, which maps observations to action distributions, maximizing the expected cumulative reward $J(\pi)=\mathop{\mathbb{E}}_{\pi}\left[\sum_{k=0}^{\infty}\gamma^{k}r_{k}\right]$, where $r_{k}=R(s_{k},a_{k})$.
@@ III-B Adversarial Inverse Reinforcement Learning
RL requires defining a reward function that accurately captures the incentive structure of real-world driving. Instead of manually defining such a reward function, we use AIRL [12] to reconstruct a surrogate reward signal from real data, given as $\mathcal{D}=\{(o_{1},a_{1}),(o_{2},a_{2}),\dots\}$. In AIRL, an additional discriminator model $D_{\phi}$ is trained to distinguish generated from real samples, producing the probability $D_{\phi}(o,a)\in\left[0,1\right]$ for the observation-action pair being real, i. e., stemming from $\mathcal{D}$. The policy is trained via RL using the surrogate reward
@@
which guides the policy toward taking realistic actions. Following [20], to smooth the decision boundary during discriminator training, we add random noise to the actions in $\mathcal{D}$, matching the standard deviation of the policy.
Figure 2: Single simulation step: The behavior model maps observations to actions, which are executed via a kinematic bicycle model for vehicles and kinematic unicycle model for pedestrians.
$t\!=\!$0\text{\,}\mathrm{s}\text{/}$\$
$t\!=\!$2\text{\,}\mathrm{s}\text{/}$\$
$t\!=\!$4\text{\,}\mathrm{s}\text{/}$\$
$t\!=\!$6\text{\,}\mathrm{s}\text{/}$\$
$t\!=\!$8\text{\,}\mathrm{s}\text{/}$\$
$t\!=\!$10\text{\,}\mathrm{s}\text{/}$\$
Figure 3: Illustration of an example scenario using instance-centric observations. Simulated vehicles and their counterparts from the real-world dataset are depicted as solid and outlined rectangles, respectively.
Figure 4: Illustration of the proposed instance-centric behavior model mapping observations to actions. Instance encoders convert observations into latent tokens. These tokens are then augmented with positional encodings relative to the target agent before being passed through multiple layers of our refinement module. Lastly, each refined agent token is decoded into its corresponding actions. Static map tokens can be reused across simulation steps.
@@
However, using a fixed $c$ across experiments, e. g., when comparing different discriminator architectures or scene representations, may result in varying expected values of $\tilde{r}$, thereby influencing the agent’s incentive to survive. For instance, a weaker discriminator tends to learn a less distinct decision boundary, yielding higher surrogate rewards, whereas a stronger discriminator learns sharper boundaries, producing lower surrogate rewards. Consequently, policies trained with stronger discriminators have less incentive to avoid termination, i. e., collisions or off-track driving, despite in both cases the optimal behavior w. r. t. the discriminator is to match the distribution of the real data $\mathcal{D}$.
To decouple the agent’s survival incentive from the discriminator’s accuracy, we propose an adaptive reward offset
@@
where $\tilde{r}_{\mathrm{mean}}(n)$ is the average surrogate reward for the generated samples in epoch $n$, and $\tilde{r}_{\mathrm{target}}$ is a hyperparameter, specifying the desired average reward. In other words, our dynamic reward offset can be interpreted as an additional reward term that encourages the policy to survive longer, especially when the discriminator is confident that the given samples are generated. This adaptive approach keeps the expected value of $\tilde{r}$ consistent, thus enabling fair comparisons across experiments, while still allowing the discriminator to guide the policy toward more realistic behavior by computing the appropriate offset to $\tilde{r}_{\mathrm{target}}$.
@@ III-C Instance-Centric Observations
For static map elements, such as pedestrian crossings or lane boundaries, we use a vectorized representation [13]. Each map element is approximated by a polyline with a maximum total length of $10\text{\,}\mathrm{m}$. A polyline consists of multiple vectors, each described by a set of features $\mathbf{v}=\left[\mathbf{v}_{\mathrm{start}},\mathbf{v}_{\mathrm{end}},\mathbf{v}_{\mathrm{type}}\right]^{\intercal}$, which includes the vector’s start and end points, as well as a one-hot encoding for the polyline type. For each polyline, the origin of its instance-centric coordinate frame is defined as the mean position of all polyline points. The x-axis is aligned with the mean direction of the vectors forming the polyline.
The local coordinate frame of an agent is determined by its last position and moving direction. Agents are described by their size (width and length), velocity, current speed limit, and a binary indicator for Vulnerable Road Users (VRUs): $\mathbf{x}=\left[\mathbf{x}_{\mathrm{size}},x_{\mathrm{vel}},x_{\mathrm{limit}},x_{\mathrm{vru}}\right]^{\intercal}$. Note that, compared to existing works utilizing agent-centric observations, here the agent’s position and orientation are not included in the agent features, but are now captured by the position and orientation of its local coordinate frame.
@@ III-D Behavior Model Details
followed by an element-wise max-pooling operation over $\{\mathbf{v}_{1},\mathbf{v}_{2},\dots,\mathbf{v}_{L}\}$ to produce the final polyline token $\mathbf{z}^{\mathrm{(map)}}$. We implement $g_{\mathrm{enc}}(\cdot)$ as a multi-layer perceptron (MLP), $f_{\mathrm{agg}}(\cdot)$ as element-wise max-pooling, and $f_{\mathrm{rel}}(\cdot)$ as concatenation. Conversely, each agent feature vector $\mathbf{x}$ is encoded into the same latent space using another MLP, producing the agent token $\mathbf{z}^{\mathrm{(agent)}}=\mathrm{MLP}(\mathbf{x})$.
Note that using agent-centric observations, requires each instance to be encoded $N_{a}\times H$ times — once per agent per simulation step, where $H$ denotes the total number of steps. In contrast, with our instance-centric representation, the polyline tokens $Z^{\mathrm{(map)}}=\{\mathbf{z}^{\mathrm{(map)}}_{i}\mid i=1,\cdots,N_{p}\}$ remain unchanged throughout the rollout. This enables reusing the tokens in subsequent simulation steps, reducing the number of encoding a polyline-instance from $N_{a}\times H$ times to just one. However, agent states evolve during the simulation rollouts, requiring agent tokens $Z^{\mathrm{(agent)}}=\{\mathbf{z}^{\mathrm{(agent)}}_{i}\mid i=1,\cdots,N_{a}\}$ to be regenerated at each simulation step. Nonetheless, since agents are encoded in a viewpoint-invariant manner, they only must be encoded $H$ times instead of $N_{a}\times H$ times.
@@
For that, we define the relative positional encoding between a target agent $i\in\{1,\cdots,N_{a}\}$ and any of the combined instance tokens $Z=\left[Z^{\mathrm{(agent)}},\;Z^{\mathrm{(map)}}\right]$, indexed by $j$, as $\mathbf{c}_{i\rightarrow j}=\mathbf{r}_{i\rightarrow j}\oplus\mathds{1}^{(\mathrm{agent})}_{j}\oplus\mathds{1}^{(\mathrm{route})}_{j}$, where $\oplus$ denotes concatenation. The binary indicators $\mathds{1}^{(\mathrm{agent})}_{j}$ and $\mathds{1}^{(\mathrm{route})}_{j}$ specify whether instance $j$ is an agent, or a map polyline that is part of the route that agent $i$ should follow, respectively. These indicators are appended here rather than to the vector features $\mathbf{v}$, as each agent may follow a different route, making a global encoding at the vector level infeasible. Finally, following the approach proposed in [24], we compute pairwise encodings by applying a feature-wise linear transformation:
@@
Decoder: The decoder processes the refined instance token $\mathbf{z}_{i}^{(K)}$ from the last Perceiver layer. In the behavior model, it outputs the mean and standard deviation of the next action. In the discriminator, which evaluates observation-action pairs, the action is appended to $\mathbf{z}_{i}^{(K)}$ to then compute the classification score $D_{\phi}$.
We use three layers of the message passing mechanism defined in (5) and $16$ channels per head in (7). Each MLP consists of a linear layer, followed by layer normalization, ReLU activation, and a second linear layer.